Massachusetts Burger is a national chain that, because of its success, has decided to expand. MassBurger, as it’s also known, is currently unlevered and has assets valued at 3750000. Suppose the size of the assets increases by a factor of 3.15 (for example, if the factor is 10, then it means the amount of assets will be 10 times as large), with all of the expansion funds coming from debt capital. What must the percentage fall in the assets be before all the equity is wiped out? Answer: % Place your answer in percentage form and two decimal places. For example, if your answer is twelve point seven five percent then submit as your answer 12.75%
Answer #1

Debt=3750000*(3.15-1)=8062500
Equity=3750000

Now, if assets fall by x%
Equity=3750000*3.15*(1-x%)-8062500
=>3750000*3.15*(1-x%)-8062500=0
=>x=31.75%